OpenStudy (anonymous):

find the absolute extrema of the function: y = 3x^(2/3) - 2x, [-1, 1]

OpenStudy (anonymous):

\[y' = 2*\frac{(1-x^{1/3})}{x^{1/3}}\] my book reads "0" as a critical number; however, after solving for both fractions = 0, i get: \[1 - x^{1/3} = 0\] \[x = 1\]

OpenStudy (anonymous):

1 should be the critical number, not zero, no?

OpenStudy (anonymous):

It looks like they both are critical values.

OpenStudy (anonymous):

can you explain how you found 0 as the critical value.

OpenStudy (anonymous):

\[\frac{2}{x^{1/3}} = 0\] \[2 = 0\] ... this is the second fraction.